OpenStudy (anonymous):

: 2(x + 6) − (x + 1) = 16 3 5 2 1

OpenStudy (theviper):

\[2x+12-x+1=16\]

OpenStudy (theviper):

\[x+13=16\]

OpenStudy (anonymous):

Maybe I'm wrong, but doesn't that +1 turn to -1 when distributing the negative sign to (x+1), Viper?

OpenStudy (theviper):

\[\LARGE{\color{red}{x=16-13=3}}\]

OpenStudy (anonymous):

2x +12 -x - 1 = 16 x +11 = 16 x = 5

OpenStudy (theviper):

Yes McGray is right

OpenStudy (theviper):

-(x+1)=-x-1